Dania is dragging a 110 kg trunk across a rough floor at a constant velocity of 5.6 m/s [E] with a horizontal force of 380 N [E].

  1. Calculate the coefficient of kinetic friction between the trunk and floor.
  2. Sonya decides to help by pulling on the trunk with a force of 150 N [Up].  Will Sonya be helping Dania?  Explain
  3. Calculate the force Dania needs to pull at now in order to keep a constant speed.
  4. Another student (not as nice as Sonya) sits on the trunk.  They have a mass of 55 kg.  What force does Dania need to pull at now for a constant speed?
1
Expert's answer
2021-10-31T17:37:14-0400

1)

"\mu=\frac F{mg}=0.35,"

2)

"\mu" will lower, she will help,

3)

"F'=F\frac {\mu}{\mu'}=325~N,"

4)

"F''=F\frac {F'}{F'+m'g}=143~N."